# Kronecker-product random matrices and a matrix least squares problem

### v2

@@ 1. Introduction
$(\mathbf{u}\otimes\mathbf{v})^{*}[G-G_{0}](\mathbf{u}^{\prime}\otimes\mathbf{v}^{\prime})\prec n^{-1/2}\text{ for deterministic unit vectors }\mathbf{u},\mathbf{v},\mathbf{u}^{\prime},\mathbf{v}^{\prime}\in\mathbb{C}^{n},$ (7)
@@
A special case of the optimization problem (1) with $\Theta=\Xi=\eta I_{n\times n}$ (and correlated Wigner inputs $A$ and $-B$) was previously analyzed in [31], in the context of the graph matching application. We review a central idea of this previous analysis in Appendix A, which, however, is special to a commutative setting and does not extend more generally. Here, we instead follow an alternative approach of a two-stage Schur-complement analysis of the resolvent, in each stage applying, in an operator-algebra setting, ideas around Dyson fixed-point equations and fluctuation averaging techniques developed for Wigner-type models in [27, 28, 24, 42] and for $\mathbb{C}^{k\times k}\otimes\mathbb{C}^{n\times n}$-valued Kronecker matrix models with fixed dimension $k$ in [5, 26]. The application of these methods in our setting of $k(n)=n$ seems to require new ideas, even to obtain optimal quantitative estimates for fixed spectral parameters on the global scale, and we discuss this further in Section 1.1 below.
@@ 1.1. Proof ideas
[…]
@@
for the linear operator $\mathcal{L}_{2}(X)=X-\frac{1}{n}\sum_{i}M_{i}XM_{i}^{*}$, with $M_{i}$ defined in (15). This type of operator associated to Dyson fixed-point equations has been studied previously in [2, 5], and we adapt a Perron-Frobenius argument of [2] to show quantitative invertibility of $\mathcal{L}_{2}$ for any fixed $z\in\mathbb{C}^{+}$. In our context, we require invertibility of $\mathcal{L}_{2}$ in the $L^{p}$-norm for each $p\in[1,\infty)$, and in infinite-dimensional settings where $\mathcal{L}_{2}$ may not have an exact Perron-Frobenius eigenvector — we thus carry out this analysis in the $L^{1}$-$L^{\infty}$ duality rather than the Hilbert-space setting of [2], and appeal to the Riesz-Thorin interpolation to obtain invertibility for all $p$ (Lemma 4.7). This shows $n^{-1}\sum_{r}G_{rj}^{(i)}G_{rj}^{(i)*}\prec n^{-1}+n^{-3\alpha}$, which applied to (14) shows the implication
@@ 2.1. Kronecker deformed Wigner model
[…]
@@ 2.2. Application to least-squares problem
Our main results for this optimization problem (30) are the following asymptotic characterization of the minimum objective value $f(\widehat{X})$ and the values of linear forms $n^{-1}{\mathbf{v}^{\prime}}^{*}\widehat{X}\mathbf{u}^{\prime}$ for deterministic test vectors $\mathbf{u}^{\prime},\mathbf{v}^{\prime}\in\mathbb{R}^{n}$.
@@
where $(\tau^{\mathcal{D}}\otimes\tau^{\mathcal{D}})[(\mathsf{a}\otimes 1+1\otimes\mathsf{b})^{2}+\Theta\otimes\Xi]^{-1}\in\mathbb{R}^{n^{2}\times n^{2}}$ is a deterministic-equivalent approximation for the matrix $[(A\otimes I+I\otimes B)^{2}+\Theta\otimes\Xi]^{-1}$ arising in the vectorization of the objective (29), and these expressions (31) and (32) coincide because this deterministic-equivalent matrix is diagonal.
In an asymptotic setting, if the empirical distributions of coordinates of $\Theta,\Xi,\mathbf{u},\mathbf{v},\mathbf{u}^{\prime},\mathbf{v}^{\prime}$ converge to deterministic limits, then a qualitative implication of Theorem 2.5 is a characterization of the almost-sure limit values of $f(\widehat{X})$ and $n^{-1}{\mathbf{v}^{\prime}}^{*}\widehat{X}\mathbf{u}^{\prime}$, as summarized in the following corollary.
@@ 2.2.1. Numerical computation
Figure 2. Values of $1/(2f(\widehat{X}))$ obtained from solving (30) across 10 independent realizations (solid dots, with vertical lines indicating 1 standard deviation) versus the theoretical prediction $T(\mathcal{P},\mathcal{Q})$ computed from Proposition 2.7 with $M=12$ (dashed lines). Here $A,B$ are GOE matrices of size $n=1000$, we take $\theta_{i}^{-1},\xi_{\alpha}^{-1},u_{i}^{2},v_{\alpha}^{2}\sim\text{Uniform}(0.05,0.5)/k$ with $k\in\{1,1.3,3\}$, and the horizontal axis indicates the correlation between coordinate pairs $(\theta_{i}^{-1},u_{i}^{2})$ and between $(\xi_{\alpha}^{-1},v_{\alpha}^{2})$.
@@ 3.1. Stochastic domination
The argument is similar to the scalar setting (c.f. [29, Lemma D.2]): For each fixed $p\in[1,\infty)$, by the triangle inequality and Hölder’s inequality for the $L^{p}$-norm (Lemma D.2),
@@ 3.3. Fluctuation averaging
The following fluctuation averaging statements in the $L^{p}$-norms on $\mathcal{X}$ are similar to those of the scalar setting, see e.g. [24, Theorem 4.7] and [30, Lemma A.2].
@@ 4.1. Operator fixed-point equations
=$\mathcal{B}^{\prime}=\{\mathsf{a}\in\mathcal{B}:\Re\mathsf{a}=\tfrac{1}{2}(\mathsf{a}+\mathsf{a}^{*})\geq\epsilon\text{ for some }\epsilon>0\}.$
@@
We adapt this idea to address two additional challenges in our setting: First, as the map $\mathsf{a}\mapsto(\mathsf{r}_{0})_{ii}\mathsf{x}\mathsf{a}\mathsf{x}(\mathsf{r}_{0})_{ii}^{*}$ is non-compact and may have continuous spectrum, we can only guarantee the existence of an approximate Perron-Frobenius eigenvector (c.f. Lemma 4.6 below). For reasons that will be clear in the proof, the approximation error must be controlled in $L^{1}$ rather than $L^{2}$, and thus we implement a version of this argument in the $L^{1}$-$L^{\infty}$ duality rather than in a Hilbert space setting, to obtain a bound for $\|\mathcal{L}_{2}^{-1}\|_{L^{\infty}\to L^{\infty}}$. Second, as we will also require a bound on $\|\mathcal{L}_{2}^{-1}\|_{L^{p}\to L^{p}}$ for each $p\in[1,\infty)$, we carry out a dual version of this argument to obtain also a bound for $\|\mathcal{L}_{2}^{-1}\|_{L^{1}\to L^{1}}$, and hence deduce a bound on $\|\mathcal{L}_{2}^{-1}\|_{L^{p}\to L^{p}}$ via the Riesz-Thorin interpolation.
@@
$01-r(\mathcal{F})\geq\delta c(\gamma,\upsilon,\delta)>$ 0 . (63)
@@
$(\mathrm{Id}-\mathcal{G}^{\prime})[\mathsf{x}(\Im\mathsf{m}_{0})\mathsf{x}]=\mathsf{w},\qquad\text{ for }\mathsf{w}=\frac{1}{n}\sum_{i=1}^{n}\mathsf{x}(\mathsf{r}_{0})_{ii}^{*}[\Im\mathsf{z}_{i}](\mathsf{r}_{0})_{ii}\mathsf{x}.$ (66)
In place of (61) and (62), we may apply $\phi(\mathsf{v}\mathsf{x}[\Im\mathsf{m}_{0}]\mathsf{x})\leq\gamma^{2}\delta^{-1}$ and $\mathsf{w}\geq\gamma^{-2}c(\gamma,\upsilon,\delta)$, where the first inequality uses $\|\mathsf{x}\|_{\mathrm{op}}\leq\gamma$ and the second uses $\|\mathsf{x}^{-1}\|_{\mathrm{op}}\leq\gamma$ and $c(\gamma,\upsilon,\delta)$ as defined in (62). Then, multiplying (66) by $\mathsf{v}$, taking the trace, and then taking the limit $\epsilon\to 0$, we obtain similarly to (63) that $01-r(\mathcal{G})\geq\delta\gamma^{-4}c(\gamma,\upsilon,\delta)>$ 0. This implies that $\mathrm{Id}-\mathcal{G}^{\prime}:L^{\infty}\to L^{\infty}$ is invertible with positivity-preserving inverse, and repeating the preceding arguments gives, for any $\mathsf{a}\in L^{\infty}$,
@@ 4.2. Weak estimates in operator norm
$\leq\frac{|r_{1}-r_{0}|}{|r_{2}-r_{0}|}\sup_{z\in S_{r_{0}}}\log|f_{i,\xi,\zeta}^{(S)}\circ M^{(iS)}(z)|+\frac{|r_{1}-r_{2}|}{|r_{2}-r_{0}|}\sup_{z\in S_{r_{2}}}\log|f_{i,\xi,\zeta}^{(S)}\circ M^{(iS)}(z)|$
@@
Note that $w(D^{(iS)})$ is independent of $\mathbf{h}_{i}^{(iS)}$, and |$|\{S\subset\{1,\ldots,n\}:|S|\leq l\}|\leq n^{l}$ and $|\mathcal{W}_{t}|=(K+2)^{t}\leq n^{C\log(K+2)}$ for all $t\leq T(n)\leq C\log n$. Then, fixing $\epsilon^{\prime}=0.05$, applying (75) to each matrix $M=w(D^{(iS)})$ conditional on $H^{(iS)}$, and taking a union bound, the event
@@
$\|\Delta_{i}^{(S)}\|_{\mathrm{op}}\leq C(\gamma,\delta)n^{-\alpha}\text{ on }\mathcal{E}.$ (85)
@@ 5. Analysis of the Kronecker deformed Wigner model
[…]
@@
and abbreviate $G=G(z)$, $\tilde{\mathsf{g}}=\tilde{\mathsf{g}}(z)$, and $\mathsf{g}=\mathsf{g}(z)$. Fix any $\epsilon>0$ and consider the event =$\mathcal{E}=\{\|B\|_{\mathrm{op}}\leq 3\}$. We apply Theorem 4.2 conditional on $B$ and this event $\mathcal{E}$, with $\mathcal{X}=\mathbb{C}^{n\times n}$, $R=G$, $\mathsf{r}_{0}=\tilde{\mathsf{g}}$, $H=A$, $\mathsf{h}=\mathsf{a}$, $\mathsf{x}=I$, and
@@
[…]
@@
Throughout this proof, $G_{ij},G_{\alpha\beta}$ etc. are all matrices in $\mathbb{C}^{n\times n}$, so the stochastic domination notation $\prec$ may be understood in the operator norm sense, c.f. Remark 3.3. We recall our convention of Roman indices $i,j,k,\ldots$ for the first tensor factor and Greek indices $\alpha,\beta,\gamma,\ldots$ for the second tensor factor. In the following, we use a superscript ()($\cdot$)( ⋅ ) to denote a minor on $A$ and []delimited-[][$\cdot$][ ⋅ ] to denote a minor on $B$.
@@ 6. Analysis of least-squares problem
Simple calculus yields the explicit forms for $\hat{\mathbf{x}}$ and $f(\hat{\mathbf{x}})$ as
@@
$-i\widetilde{P}=\begin{bmatrix}\Theta\otimes\Xi&i(A\otimes I+I\otimes B)\\ i(A\otimes I+I\otimes B)&I\otimes I\end{bmatrix}^{-1}=\begin{bmatrix}[(A\otimes I+I\otimes B)^{2}+\Theta\otimes\Xi]^{-1}&*\\ *&*\end{bmatrix}$ (121)
@@
Here, in the second line, we have applied (47) of Theorem 4.2 twice as in the proof of Theorem 2.4 in Section 5, first to the second tensor factor conditional on $B$ and the event =$\mathcal{E}=\{\|B\|_{\mathrm{op}}\leq 3\}$ with $\mathcal{X}=\mathbb{C}^{2\times 2}\otimes\mathbb{C}^{n\times n}$ (the product of first and third factors),
@@
$f(\hat{\mathbf{x}})$ $=\frac{1}{2}\frac{1}{n^{-2}(\mathbf{u}\otimes\mathbf{v})^{*}P_{0}(\mathbf{u}\otimes\mathbf{v})}+O_{\prec}(n^{-1/2}),$
@@
where $\Theta,\mathsf{u},\mathsf{u}^{\prime},\Xi,\mathsf{v},\mathsf{v}^{\prime}\in\mathcal{D}\subset\mathcal{A}$ on the left are real diagonal matrices, and $\theta,\mathsf{U},\mathsf{U}^{\prime},\xi,\mathsf{V},\mathsf{V}^{\prime}\in\mathcal{A}$ on the right are commuting, self-adjoint limiting operators, free of $(\mathsf{a},\mathsf{b})$ and for which $(\theta,\mathsf{U},\mathsf{U}^{\prime})$ and $(\xi,\mathsf{V},\mathsf{V}^{\prime})$ have joint laws under $\tau$ given by $\mathcal{P}$ and $\mathcal{Q}$, respectively. The assumptions $\|\Theta\|_{\mathrm{op}},\|\Xi\|_{\mathrm{op}}\leq\upsilon$ and $\Theta,\Xi\geq\delta$ imply that the spectrum of $(\mathsf{a}\otimes 1+1\otimes\mathsf{b})^{2}+\Theta\otimes\Xi$ is contained in $[\delta^{2},16+\upsilon^{2}]$. The inverse function $x\mapsto x^{-1}$ may be approximated uniformly by polynomials on this interval, so the above convergence implies
@@
[…]
@@
$[(\mathsf{a}\otimes 1+1\otimes\mathsf{b})^{2}+\theta\otimes\xi]^{-1}=(\mathbf{e}_{1}\otimes 1\otimes 1)^{*}[-i\eta^{-1}\tilde{\mathsf{p}}](\mathbf{e}_{1}\otimes 1\otimes 1).$ (124)
@@
[…]
@@
so $\tilde{\mathsf{q}}=\tilde{\mathsf{a}}+\tilde{\mathsf{a}}^{*}+\tilde{\mathsf{b}}+\tilde{\mathsf{b}}^{*}$ is self-adjoint with spectrum contained in $[-4\eta^{-1},4\eta^{-1}]$. Hence
@@
A direct calculation shows that this identity holds also for $m=0$, upon defining $\operatorname{NC}_{2,2}(0)$ to have the single pair (,)($\emptyset$,$\emptyset$)( ∅ , ∅ ) with its value defined as in (34). Applying this back to (126) and using $i^{m}=(-1)^{m/2}$ in $C_{m}(z)$ concludes the proof. ∎

### v3

@@ 1. Introduction
[…]
@@
A special case of the optimization problem (1) with $\Theta=\Xi=\eta I_{n\times n}$ (and correlated Wigner inputs $A$ and B-B) was previously analyzed in [31], in the context of the graph matching application. We review a central idea of this previous analysis in Appendix A, which, however, is special to a commutative setting and does not extend more generally. Here, we instead follow an alternative approach of a two-stage Schur-complement analysis of the resolvent, in each stage applying, in an operator-algebra setting, ideas around Dyson fixed-point equations and fluctuation averaging techniques developed for Wigner-type models in [27, 28, 24, 43] and for $\mathbb{C}^{k\times k}\otimes\mathbb{C}^{n\times n}$-valued Kronecker matrix models with fixed dimension $k$ in [4, 26]. The application of these methods in our setting of $k(n)=n$ seems to require new ideas, even to obtain optimal quantitative estimates for fixed spectral parameters on the global scale, and we discuss this further in Section 1.1 below.
@@ 1.1. Proof ideas
$M_{B}=(\tau\otimes I)\Big[\underbrace{(\mathsf{a}\otimes I+1\otimes B+\Theta\otimes\Xi-z\,1\otimes I)^{-1}}_{=\tilde{\mathsf{g}}}\Big$] (11)
@@
for the linear operator $\mathcal{L}_{2}(X)=X-\frac{1}{n}\sum_{i}M_{i}XM_{i}^{*}$, with $M_{i}$ defined in (15). This type of operator associated to Dyson fixed-point equations has been studied previously in [2, 4], and we adapt a Perron-Frobenius argument of [2] to show quantitative invertibility of $\mathcal{L}_{2}$ for any fixed $z\in\mathbb{C}^{+}$. In our context, we require invertibility of $\mathcal{L}_{2}$ in the $L^{p}$-norm for each $p\in[1,\infty)$, and in infinite-dimensional settings where $\mathcal{L}_{2}$ may not have an exact Perron-Frobenius eigenvector — we thus carry out this analysis in the $^{1}-LL^{\infty}$ duality rather than the Hilbert-space setting of [2], and appeal to the Riesz-Thorin interpolation to obtain invertibility for all $p$ (Lemma 4.7). This shows $n^{-1}\sum_{r}G_{rj}^{(i)}G_{rj}^{(i)*}\prec n^{-1}+n^{-3\alpha}$, which applied to (14) shows the implication
@@ 2.1. Kronecker deformed Wigner model
    $\mathcal{A}$ has two semicircular elements $\mathsf{a},\mathsf{b}$ (i.e. satisfying $\tau(\mathsf{a}^{k})=\tau(\mathsf{b}^{k})=\int_{-2}^{2}x^{k}\,\frac{1}{2\pi}\sqrt{4-x^{2}}\,dx$ for each $k\geq 1$) which are free of $\mathcal{M}$ with respect to $\tau$.
@@ 2.2. Application to least-squares problem
Our main results for this optimization problem (30) are the following asymptotic characterization of the minimum objective value f(X^)f($\widehat${X}) and the values of linear forms $n^{-1}{\mathbf{v}^{\prime}}^{*}\widehat{X}\mathbf{u}^{\prime}$ for deterministic test vectors $\mathbf{u}^{\prime},\mathbf{v}^{\prime}\in\mathbb{R}^{n}$.
@@
where $(\tau^{\mathcal{D}}\otimes\tau^{\mathcal{D}})[(\mathsf{a}\otimes 1+1\otimes\mathsf{b})^{2}+\Theta\otimes\Xi]^{-1}\in\mathbb{R}^{n^{2}\times n^{2}}$ is a deterministic-equivalent approximation for the matrix $(A\otimes I+I\otimes B)^{2}+\Theta\otimes\Xi]^{-1}$ arising in the vectorization of the objective (29), and these expressions (31) and (32) coincide because this deterministic-equivalent matrix is diagonal.
In an asymptotic setting, if the empirical distributions of coordinates of $\Theta,\Xi,\mathbf{u},\mathbf{v},\mathbf{u}^{\prime},\mathbf{v}^{\prime}$ converge to deterministic limits, then a qualitative implication of Theorem 2.5 is a characterization of the almost-sure limit values of f(X^)f($\widehat${X}) and $n^{-1}{\mathbf{v}^{\prime}}^{*}\widehat{X}\mathbf{u}^{\prime}$, as summarized in the following corollary.
@@ 2.2.1. Numerical computation
Figure 2. Values of $)1/(2f(\widehat{X}$)) obtained from solving (30) across 10 independent realizations (solid dots, with vertical lines indicating 1 standard deviation) versus the theoretical prediction $T(\mathcal{P},\mathcal{Q})$ computed from Proposition 2.7 with $M=12$ (dashed lines). Here $A,B$ are GOE matrices of size $n=1000$, we take $\theta_{i}^{-1},\xi_{\alpha}^{-1},u_{i}^{2},v_{\alpha}^{2}\sim\text{Uniform}(0.05,0.5)/k$ with $k\in\{1,1.3,3\}$, and the horizontal axis indicates the correlation between coordinate pairs $(\theta_{i}^{-1},u_{i}^{2})$ and between $(\xi_{\alpha}^{-1},v_{\alpha}^{2})$.
@@ 3.1. Stochastic domination
The argument is similar to the scalar setting (c.f. [30, Lemma D.2]): For each fixed $p\in[1,\infty)$, by the triangle inequality and Hölder’s inequality for the $L^{p}$-norm (Lemma D.2),
@@ 3.3. Fluctuation averaging
The following fluctuation averaging statements in the $L^{p}$-norms on $\mathcal{X}$ are similar to those of the scalar setting, see e.g. [24, Theorem 4.7] and [29, Lemma A.2].
@@ 4.1. Operator fixed-point equations
$\mathcal{B}^{\prime}=\{\mathsf{a}\in\mathcal{B}:\Re\mathsf{a}=\tfrac{1}{2}(\mathsf{a}+\mathsf{a}^{*})\geq\epsilon\text{ for some }\epsilon>0\}.$
@@
We adapt this idea to address two additional challenges in our setting: First, as the map $\mathsf{a}\mapsto(\mathsf{r}_{0})_{ii}\mathsf{x}\mathsf{a}\mathsf{x}(\mathsf{r}_{0})_{ii}^{*}$ is non-compact and may have continuous spectrum, we can only guarantee the existence of an approximate Perron-Frobenius eigenvector (c.f. Lemma 4.6 below). For reasons that will be clear in the proof, the approximation error must be controlled in $L^{1}$ rather than $L^{2}$, and thus we implement a version of this argument in the $^{1}-LL^{\infty}$ duality rather than in a Hilbert space setting, to obtain a bound for $\|\mathcal{L}_{2}^{-1}\|_{L^{\infty}\to L^{\infty}}$. Second, as we will also require a bound on $\|\mathcal{L}_{2}^{-1}\|_{L^{p}\to L^{p}}$ for each $p\in[1,\infty)$, we carry out a dual version of this argument to obtain also a bound for $\|\mathcal{L}_{2}^{-1}\|_{L^{1}\to L^{1}}$, and hence deduce a bound on $\|\mathcal{L}_{2}^{-1}\|_{L^{p}\to L^{p}}$ via the Riesz-Thorin interpolation.
@@
$1-r(\mathcal{F})\geq\delta c(\gamma,\upsilon,\delta)>0.$ (63)
@@
[…]
In place of (61) and (62), we may apply $\phi(\mathsf{v}\mathsf{x}[\Im\mathsf{m}_{0}]\mathsf{x})\leq\gamma^{2}\delta^{-1}$ and $\mathsf{w}\geq\gamma^{-2}c(\gamma,\upsilon,\delta)$, where the first inequality uses $\|\mathsf{x}\|_{\mathrm{op}}\leq\gamma$ and the second uses $\|\mathsf{x}^{-1}\|_{\mathrm{op}}\leq\gamma$ and $c(\gamma,\upsilon,\delta)$ as defined in (62). Then, multiplying (66) by $\mathsf{v}$, taking the trace, and then taking the limit $\epsilon\to 0$, we obtain similarly to (63) that $1-r(\mathcal{G})\geq\delta\gamma^{-4}c(\gamma,\upsilon,\delta)>0$. This implies that $\mathrm{Id}-\mathcal{G}^{\prime}:L^{\infty}\to L^{\infty}$ is invertible with positivity-preserving inverse, and repeating the preceding arguments gives, for any $\mathsf{a}\in L^{\infty}$,
@@ 4.2. Weak estimates in operator norm
$\leq\frac{|r_{1}-r_{2}|}{|r_{2}-r_{0}|}\sup_{z\in S_{r_{0}}}\log|f_{i,\xi,\zeta}^{(S)}\circ M^{(iS)}(z)|+\frac{|r_{1}-r_{0}|}{|r_{2}-r_{0}|}\sup_{z\in S_{r_{2}}}\log|f_{i,\xi,\zeta}^{(S)}\circ M^{(iS)}(z)|$
@@
Note that $w(D^{(iS)})$ is independent of $\mathbf{h}_{i}^{(iS)}$, and $|\{S\subset\{1,\ldots,n\}:|S|\leq l\}|\leq n^{l}$ and $|\mathcal{W}_{t}|=(K+2)^{t}\leq n^{C\log(K+2)}$ for all $t\leq T(n)\leq C\log n$. Then, fixing $\epsilon^{\prime}=0.05$, applying (75) to each matrix $M=w(D^{(iS)})$ conditional on $H^{(iS)}$, and taking a union bound, the event
@@
[…]
@@ 5. Analysis of the Kronecker deformed Wigner model
$\mathsf{g}(z)$ $=(\mathsf{q}-z\,1\otimes 1)^{-1},\qquad G_{0}(z)=(\tau^{\mathcal{D}}\otimes\tau^{\mathcal{D}})[\mathsf{g}(z)],\qquad m_{0}(z)=\tau\otimes\tau[\mathsf{g}(z)].$
@@
and abbreviate $G=G(z)$, $\tilde{\mathsf{g}}=\tilde{\mathsf{g}}(z)$, and $\mathsf{g}=\mathsf{g}(z)$. Fix any $\epsilon>0$ and consider the event $\mathcal{E}=\{\|B\|_{\mathrm{op}}\leq 3\}$. We apply Theorem 4.2 conditional on $B$ and this event $\mathcal{E}$, with $\mathcal{X}=\mathbb{C}^{n\times n}$, $R=G$, $\mathsf{r}_{0}=\tilde{\mathsf{g}}$, $H=A$, $\mathsf{h}=\mathsf{a}$, $\mathsf{x}=I$, and
@@
$\mathsf{z}={-}\mathsf{a}\otimes I-\Theta\otimes\Xi+z\,1_{\mathcal{A}}\otimes I.$
@@
Throughout this proof, $G_{ij},G_{\alpha\beta}$ etc. are all matrices in $\mathbb{C}^{n\times n}$, so the stochastic domination notation $\prec$ may be understood in the operator norm sense, c.f. Remark 3.3. We recall our convention of Roman indices $i,j,k,\ldots$ for the first tensor factor and Greek indices $\alpha,\beta,\gamma,\ldots$ for the second tensor factor. In the following, we use a superscript $(\cdot)$ to denote a minor on $A$ and $\cdot$] to denote a minor on $B$.
@@ 6. Analysis of least-squares problem
[…]
@@
$i\widetilde{P}=\begin{bmatrix}\Theta\otimes\Xi&i(A\otimes I+I\otimes B)\\ i(A\otimes I+I\otimes B)&I\otimes I\end{bmatrix}^{-1}=\begin{bmatrix}[(A\otimes I+I\otimes B)^{2}+\Theta\otimes\Xi]^{-1}&*\\ *&*\end{bmatrix}$ (121)
@@
Here, in the second line, we have applied (47) of Theorem 4.2 twice as in the proof of Theorem 2.4 in Section 5, first to the second tensor factor conditional on $B$ and the event $\mathcal{E}=\{\|B\|_{\mathrm{op}}\leq 3\}$ with $\mathcal{X}=\mathbb{C}^{2\times 2}\otimes\mathbb{C}^{n\times n}$ (the product of first and third factors),
@@
[…]
@@
where $\Theta,\mathsf{u},\mathsf{u}^{\prime},\Xi,\mathsf{v},\mathsf{v}^{\prime}\in\mathcal{D}\subset\mathcal{A}$ on the left are real diagonal matrices, and $\theta,\mathsf{U},\mathsf{U}^{\prime},\xi,\mathsf{V},\mathsf{V}^{\prime}\in\mathcal{A}$ on the right are commuting, self-adjoint limiting operators, free of $(\mathsf{a},\mathsf{b})$ and for which $(\theta,\mathsf{U},\mathsf{U}^{\prime})$ and $(\xi,\mathsf{V},\mathsf{V}^{\prime})$ have joint laws under $\tau$ given by $\mathcal{P}$ and $\mathcal{Q}$, respectively. The assumptions $\|\Theta\|_{\mathrm{op}},\|\Xi\|_{\mathrm{op}}\leq\upsilon$ and $\Theta,\Xi\geq\delta$ imply that the spectrum of $(\mathsf{a}\otimes 1+1\otimes\mathsf{b})^{2}+\Theta\otimes\Xi$ is contained in $\delta^{2},16+\upsilon^{2}]$. The inverse function $x\mapsto x^{-1}$ may be approximated uniformly by polynomials on this interval, so the above convergence implies
@@
$\tilde{\mathsf{p}}=\begin{bmatrix}-i\eta^{-1}\,\theta\otimes\xi&\mathsf{a}\otimes 1+1\otimes\mathsf{b}\\ \mathsf{a}\otimes 1+1\otimes\mathsf{b}&-i\eta\,1\otimes 1\end{bmatrix}^{-1}.$
@@
$(\mathsf{a}\otimes 1+1\otimes\mathsf{b})^{2}+\theta\otimes\xi]^{-1}=(\mathbf{e}_{1}\otimes 1\otimes 1)^{*}[-i\eta^{-1}\tilde{\mathsf{p}}](\mathbf{e}_{1}\otimes 1\otimes 1).$ (124)
@@
$\tilde{\mathsf{p}}$ $=\left(\tilde{\mathsf{d}}^{-1}\begin{bmatrix}-i\,1\otimes 1&\mathsf{d}_{a}\mathsf{a}\otimes\mathsf{d}_{b}+\mathsf{d}_{a}\otimes\mathsf{d}_{b}\mathsf{b}\\ \mathsf{a}\mathsf{d}_{a}\otimes\mathsf{d}_{b}+\mathsf{d}_{a}\otimes\mathsf{b}\mathsf{d}_{b}&-i\,1\otimes 1\end{bmatrix}\tilde{\mathsf{d}}^{-1}\right)^{-1}$
@@
so $\tilde{\mathsf{q}}=\tilde{\mathsf{a}}+\tilde{\mathsf{a}}^{*}+\tilde{\mathsf{b}}+\tilde{\mathsf{b}}^{*}$ is self-adjoint with spectrum contained in $4\eta^{-1},4\eta^{-1}]$. Hence
@@
A direct calculation shows that this identity holds also for $m=0$, upon defining $\operatorname{NC}_{2,2}(0)$ to have the single pair $(\emptyset,\emptyset)$ with its value defined as in (34). Applying this back to (126) and using $i^{m}=(-1)^{m/2}$ in $C_{m}(z)$ concludes the proof. ∎